Kimetics and Equlibtiunm 6. For reaction 1, the K, equals 68 × 10s at 298 K. N,(g) + 3H2(g) 근 2 NH,(g) Answer the following questions in alphabetical order to determine the K, for reaction 2 NH,() 1/2 N,(g) + 3/2H,(g) (2) a. What is the equilibtrium constant expres sion for reaction 1 as a K expression? b. What is the equilibrium constant expres- sion for reaction 1 as a K, expression? c. What is the equilbium constant expres- sion for reaction 2 as a K, expression? d. What is the equilbrium constant expres- sion for reaction 2 as a K expression? e. How are reactions 1 and 2 related? f. How are the equlibrium constant expressions related for reactions 1 and 2 g What is the equation used to relate K, and K? h. What is án for reaction 1? For teactoon 2? k. What is the K, for reaction 1 i. What is the K, for reaction 2 j. What is the K, for reaction 2? 1 What is the K, for reaction 2?

kinetics and equilibrium 6.
Answer #1

(a)

Kp = PNH32 / PN2*PH23

(b)

Kc = [NH3]2 / [N2][H2]3

(c)

Kp = PN21/2 * PH23/2 / PNH3

(d)

Kc = [N2]1/2 * [H2]3/2 / [NH3]

(e)

Reaction 1 is the formation of 2 mol of Ammonia

Where as reaction 2 is the decomposition of 1 mol of ammonia

(f)

(g)

Kp = Kc(RT)\Deltan

(h)

For reaction (1)

\Deltan = 2 - ( 1 + 3 ) = - 2

For reaction (2)

\Deltan = (1/2) + (3/2) - 1 = 1

(i)

Kp2 = 1 / sqrt.Kp1 = 1 / sqrt.(6.8 * 105)

Kp2 = 1.21 * 10-3

(j)

Kp2 = Kc2 * (RT)\Deltan

Kc2 = 1.21 * 10-3 / (8.314 * 298)1

Kc2 = 4.89 * 10-7

(k)

Kp1 = Kc1 * (RT)\Deltan

6.8 * 105 = Kc1 * (8.314 * 298)-2

Kc1 = 4.17 * 1012